# Optimal low-rank approximations for linear Gaussian inverse problems on Hilbert spaces Part II: posterior mean approximation

### v2

@@ 5 Optimal approximation of the mean
Similarly to [33, Section 4.1], a comparison between the minimal losses of Theorem 5.10 and Theorem 5.11 gives us insight as to which approximation procedure is preferable in a specific setting. As the theorems show, the decay of the eigenvalues $(\lambda_{i})_{i}$ of $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$ governs this choice. The loss of the optimal approximation in Theorem 5.10 and in Theorem 5.11 is $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})$ and $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{3}$ respectively. If $\tfrac{-\lambda_{i}}{1+\lambda_{i}}\leq 1$ or equivalently $-\lambda_{i}\leq\tfrac{1}{2}$ for every $i>r$, then we have $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})\geq\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{3}$. Since the sequence $(\lambda_{i})_{i}\subset(-1,0]$ increases to zero by Proposition 3.4, and since $(\lambda_{i})_{i}$ have the interpretation of variance reduction by the discussion after Proposition 3.6, it follows that if there exists some $r<n$ such that the relative variance reduction along $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ is smaller than $\tfrac{1}{2}$ for $i>r$, then the loss $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{3}$ that arises from exploiting the structure (3a) of the posterior mean is smaller than the loss that ignores this structure. In other words, one can achieve on average a smaller loss in the posterior mean approximation that exploits the structure (3a) of the posterior mean, if the ratio of the posterior variance to the prior variance along $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ decays below the threshold of $\tfrac{1}{2}$ for sufficiently large $i$. If for example $\lambda_{i}>-\tfrac{1}{2}$ for every $i\in\mathbb{N}$, then this decay does not occur, and one can obtain a smaller loss by ignoring the structure.
@@ 6 Optimal joint approximation of the mean and covariance
In Section 4, we considered the optimal rank-$r$ approximation of the posterior covariance given the same mean, while in Section 5 we considered the optimal rank-$r$ approximation of the posterior mean given the same posterior covariance. In this section, we consider jointly approximating the posterior mean and covariance in the reverse KL divergence defined in Section 3. Approximation in reverse KL divergence is important in the context of variational inference, c.f. [27, Theorem 5]. We leave the solution of the optimal joint approximation of the mean and covariance for the forward KL divergence for future work.
@@
which constitutes a Pythagorean-like identity for the Kullback–Leibler divergence between two Gaussians. The identity above is reasonable, since the Kullback–Leibler divergence is a Bregman divergence, which are known to satisfy generalised Pythagorean theorems. See e.g. [1, Section 1.6] or [24] for the information geometry perspective on Pythagorean identities and [22, Theorem 2.1] for a Pythagorean theorem in the context of dimension reduction for Bayesian inverse problems.
@@ 7 Characterisation through optimal projection
where again $X\sim\mu_{\textup{pr}}=\mathcal{N}(0,\mathcal{C}_{\textup{pr}})$. This problem only differs from Section 2 in the replacement of the forward map $G$ by $GP$. As before, we denote by $y$ an arbitrary realisation of $Y$. Let $\mu_{P,\textup{pos}}(y)\sim\mathcal{N}(m_{P,\textup{pos}}(y),\mathcal{C}_{P,\textup{pos}})$ be the posterior distribution corresponding to (29) and $\mu_{\textup{pr}}=\mathcal{N}(0,\mathcal{C}_{\textup{pr}})$. Because $GP$ is continuous, it follows from [34, Theorem 6.31] that $\mu_{P,\textup{pos}}(y)\sim\mu_{\textup{pr}}\sim\mu_{\textup{pos}}(y)$, where $\mu_{\textup{pos}}(y)$ is the posterior distribution of the full observation model (1). For the chosen value of $r$ and $i=1,2$, let $\hat{\mu}^{(i)}_{\textup{pos}}(y)=\mathcal{N}(\hat{m}^{(i)}_{\textup{pos}}(y),\mathcal{C}^{\textup{opt}}_{r})$ denote the on average optimal posterior approximation of $\mu_{\textup{pos}}(y)$ obtained in Section 6. Thus, $\mathcal{C}^{\textup{opt}}_{r}$ is given by Theorem 4.2 and $\hat{m}^{(i)}_{\textup{pos}}(y)=A^{\textup{opt},(i)}_{r}y$ is given by Theorem 5.11 for $i=1$ and Theorem 5.10 for $i=2$. Proposition 6.1 implies that $\mathbb{E}\left[D_{\textup{KL}}(\mu_{P,\textup{pos}}(Y)\|\mu_{\textup{pos}}(Y))\right]\geq\mathbb{E}\left[D_{\textup{KL}}(\hat{\mu}_{\textup{pos}}(Y)\|\mu_{\textup{pos}}(Y))\right]$. For $i=2$, we show that this lower bound is attained, that is, there exists a suitable choice of $P$ such that for every realisation $y$ we have $\mu_{P,\textup{pos}}(y)=\hat{\mu}_{\textup{pos}}(y)$. The proof is given in Section B.3.
@@
In the finite-dimensional setting, it is shown in [33, Corollary 3.2] that the posterior covariance corresponding to the model (29) agrees with the solution of Problem 4.1. Proposition 7.1 generalises this to infinite dimensions and adds an analogous statement for the posterior mean of model (29): the exact posterior mean of the projected problem (29) with $P$ as in Proposition 7.1 is equal to the optimal low-rank structure-ignoring posterior mean approximation given by Theorem 5.10.
@@ 8 Examples
In this section we consider two examples of linear Gaussian inverse problems, and identify the prior-preconditioned Hessian (9a) and its non-self adjoint square root (20) in terms of the functions occurring in the forward problem and the prior. Further details and explanations are provided in Appendix C.
@@
$\partial_{t}u-\partial_{xx}u$ $=0,$ in ($\text{in }(0,1)\times(0,T),$
[…]
$u(0,\cdot)=u(1,\cdot)$ $=0,$ on ($\text{on }(0,T].$
The true initial state $x^{\dagger}$ is unknown and needs to be estimated from noisy observations of $u$ at $(x_{i},t_{i})_{i=1}^{n}\subset(0,1)\times(0,T]$. We assume i.i.d. standard Gaussian noise. This problem is similar to [34, Example 3.5] and [15, Section 4.2]. However, in this example we do not observe the entire spatial temperature profile, but observe at finitely many fixed spatial locations, and we consider periodic instead of Dirichlet boundary conditions.
@@ 9 Conclusion
In this work, we have provided the formulation and solution of the low-rank posterior mean approximation problem directly on infinite-dimensional separable Hilbert spaces. We considered approximations that ignore and preserve the structure of the prior-to-posterior mean update in Theorem 5.10 and Theorem 5.11 respectively. To quantify the posterior mean approximation quality, we have considered various loss classes. These loss classes consist of divergences between the exact Gaussian posterior and the approximate Gaussian posterior given by an approximate posterior mean and the exact posterior covariance, after averaging over the data distribution. The chosen divergences are the Hellinger distance and the Renyi, Amari, and forward and reverse KL divergences. These loss classes form a natural extension of the Bayes risk used in finite dimensions in [33], and were used to assess optimality for low-rank approximations to the posterior covariance in Part I of this work, see [6].
@@ A.1 Hilbert spaces and bounded operators
Suppose that $(h_{n})_{n}\subset\mathcal{H}$ is weakly convergent with limit $h\in\mathcal{H}$, i.e. $\langle h_{n},k\rangle\rightarrow\langle h,k\rangle$ for all $k\in\mathcal{H}$ as $n\rightarrow\infty$. In particular, $\langle Ah,h-h_{n}\rangle\rightarrow 0$. Since the sequence $(\langle h_{n},k\rangle)_{n}$ is bounded for each $k\in\mathcal{H}$, the principle of uniform boundedness, c.f. [7, Theorem III.14.3], implies that $(h_{n})_{n}$ is a bounded sequence. By [28, Theorem VI.11], $(Ah_{n})_{n}$ converges in norm to $Ah$ since $A$ is compact. Thus, $\lvert\langle A(h-h_{n}),h_{n}\rangle\rvert\leq\lVert A(h-h_{n})\rVert\sup_{n}\lVert h_{n}\rVert\rightarrow 0$. We conclude that $\lvert\langle Ah,h\rangle-\langle Ah_{n},h_{n}\rangle\rvert\leq\lvert\langle A(h-h_{n}),h_{n}\rangle\rvert+\lvert\langle Ah,h-h_{n}\rangle\rvert\rightarrow 0$. ∎
@@ A.2 Unbounded operators
$\operatorname{dom}{A^{*}}\coloneqq\{k\in\mathcal{K}:\ h\mapsto\langle Ah,k\rangle\text{ is a bounded linear functional on }\operatorname{dom}{A}\}.$
@@ B.1 Proofs of Section 3
By Lemma A.11(i), $(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\in\mathcal{B}(\mathcal{H})$ satisfies $(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}w_{i}=\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$. The above display thus shows $I-\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}=\sum_{i}-\lambda_{i}w_{i}\otimes w_{i}$. This is a nonnegative and compact operator, since $(-\lambda_{i})_{i}\in\ell^{2}([0,1))$. Applying [18, eq. (4.13)] to this operator, we get for any subspace $V_{r}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ of dimension $r$,
@@
where in the final step we use that the ratio $\frac{\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,z\rangle)}{\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,z\rangle)}$ is invariant under scaling of $X$. It remains to show that the final infimum above is attained. Since $\{z\in\mathcal{H}:\ \lVert z\rVert\leq 1\}$ is weakly compact by [7, Theorem V.4.2], the closed subspace $(\mathcal{C}_{\textup{pr}}^{-1/2}V_{r})^{\perp}\cap\{z\in\mathcal{H}:\ \lVert z\rVert=1\}$ of $\{z\in\mathcal{H}:\ \lVert z\rVert\leq 1\}$ is also weakly compact. Furthermore, $\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,z\rangle)=\langle\mathcal{C}_{\textup{pos}}z,z\rangle$ by Lemma A.15, which is weakly continuous in $z$ by Lemma A.9. Similarly, $\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,z\rangle)$ is weakly continuous. Thus the ratio $\frac{\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,z\rangle)}{\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,z\rangle)}$ is weakly continuous on the weakly compact set $(\mathcal{C}_{\textup{pr}}^{-1/2}V_{r})^{\top}\cap\{z\in\mathcal{H}:\ \lVert z\rVert=1\}$. It follows that the infima above are attained, proving (11). ∎
@@ B.2 Proofs of Section 5
This shows that $\mathcal{C}_{\textup{pos}}=S_{\textup{pos}}S_{\textup{pos}}^{*}$, which proves item (i). Item (ii) now immediately follows from [12, Corollary B.3] and the equality $\operatorname{ran}{S_{\textup{pos}}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}.$ For item (iii), we note that by (22a) we have for $h\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, $\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}h=\sum_{i}(1+\lambda_{i})^{1/2}\langle h,w_{i}\rangle w_{i}=h-\sum_{i=1}^{n}\langle h,w_{i}\rangle w_{i}+\sum_{i=1}^{n}(1+\lambda_{i})^{1/2}\langle h,w_{i}\rangle w_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ as a sum of elements of $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, because $(w_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ by Proposition 3.4. Furthermore, if $k\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, then $h\coloneqq\sum_{i}(1+\lambda_{i})^{-1/2}\langle k,w_{i}\rangle w_{i}$ satisfies $h=k-\sum_{i=1}^{n}\langle k,w_{i}\rangle w_{i}+\sum_{i=1}^{n}(1+\lambda_{i})^{-1/2}\langle k,w_{i}\rangle w_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. By (22a), we have $\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}h=\sum_{i}(1+\lambda_{i})^{1/2}\langle h,w_{i}\rangle w_{i}=\sum_{i}\langle k,w_{i}\rangle w_{i}=k$. We conclude that $\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}$ maps $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ onto $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, so that

### v3

@@ 5 Optimal approximation of the mean
Similarly to [39, Section 4.1], a comparison between the minimal losses of Theorem 5.10 and Theorem 5.11 gives us insight as to which approximation procedure is preferable in a specific setting. As the theorems show, the decay of the eigenvalues $(\lambda_{i})_{i}$ of $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$ governs this choice. The loss of the optimal approximation in Theorem 5.10 and in Theorem 5.11 is $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})$ and $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{3}$ respectively. If $\tfrac{-\lambda_{i}}{1+\lambda_{i}}\leq 1$ or equivalently $\lambda_{i}\leq\tfrac{1}{2}$ for every $i>r$, then we have $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})\geq\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{3}$. Since the sequence $(\lambda_{i})_{i}\subset(-1,0]$ increases to zero by Proposition 3.4, and since $(\lambda_{i})_{i}$ have the interpretation of variance reduction by the discussion after Proposition 3.6, it follows that if there exists some $r<n$ such that the relative variance reduction along $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ is smaller than $\tfrac{1}{2}$ for $i>r$, then the loss $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{3}$ that arises from exploiting the structure (3a) of the posterior mean is smaller than the loss that ignores this structure. In other words, one can achieve on average a smaller loss in the posterior mean approximation that exploits the structure (3a) of the posterior mean, if the ratio of the posterior variance to the prior variance along $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ decays below the threshold of $\tfrac{1}{2}$ for sufficiently large $i$. If for example $\lambda_{i}>-\tfrac{1}{2}$ for every $i\in\mathbb{N}$, then this decay does not occur, and one can obtain a smaller loss by ignoring the structure.
@@ 6 Optimal joint approximation of the mean and covariance
In Section 4, we considered the optimal rank-$r$ approximation of the posterior covariance given the same mean, while in Section 5 we considered the optimal rank-$r$ approximation of the posterior mean given the same posterior covariance. In this section, we consider jointly approximating the posterior mean and covariance in the reverse KL divergence defined in Section 3. Approximation in reverse KL divergence is important in the context of variational inference, c.f. [32, Theorem 5]. We leave the solution of the optimal joint approximation of the mean and covariance for the forward KL divergence for future work.
@@
which constitutes a Pythagorean-like identity for the Kullback–Leibler divergence between two Gaussians. The identity above is reasonable, since the Kullback–Leibler divergence is a Bregman divergence, which are known to satisfy generalised Pythagorean theorems. See e.g. [1, Section 1.6] or [29] for the information geometry perspective on Pythagorean identities and [27, Theorem 2.1] for a Pythagorean theorem in the context of dimension reduction for Bayesian inverse problems.
@@ 7 Characterisation through optimal projection
where again $X\sim\mu_{\textup{pr}}=\mathcal{N}(0,\mathcal{C}_{\textup{pr}})$. This problem only differs from Section 2 in the replacement of the forward map $G$ by $GP$. As before, we denote by $y$ an arbitrary realisation of $Y$. Let $\mu_{P,\textup{pos}}(y)\sim\mathcal{N}(m_{P,\textup{pos}}(y),\mathcal{C}_{P,\textup{pos}})$ be the posterior distribution corresponding to (29) and $\mu_{\textup{pr}}=\mathcal{N}(0,\mathcal{C}_{\textup{pr}})$. Because $GP$ is continuous, it follows from [40, Theorem 6.31] that $\mu_{P,\textup{pos}}(y)\sim\mu_{\textup{pr}}\sim\mu_{\textup{pos}}(y)$, where $\mu_{\textup{pos}}(y)$ is the posterior distribution of the full observation model (1). For the chosen value of $r$ and $i=1,2$, let $\hat{\mu}^{(i)}_{\textup{pos}}(y)=\mathcal{N}(\hat{m}^{(i)}_{\textup{pos}}(y),\mathcal{C}^{\textup{opt}}_{r})$ denote the on average optimal posterior approximation of $\mu_{\textup{pos}}(y)$ obtained in Section 6. Thus, $\mathcal{C}^{\textup{opt}}_{r}$ is given by Theorem 4.2 and $\hat{m}^{(i)}_{\textup{pos}}(y)=A^{\textup{opt},(i)}_{r}y$ is given by Theorem 5.11 for $i=1$ and Theorem 5.10 for $i=2$. Proposition 6.1 implies that $\mathbb{E}\left[D_{\textup{KL}}(\mu_{P,\textup{pos}}(Y)\|\mu_{\textup{pos}}(Y))\right]\geq\mathbb{E}\left[D_{\textup{KL}}(\hat{\mu}_{\textup{pos}}(Y)\|\mu_{\textup{pos}}(Y))\right]$. For $i=2$, we show that this lower bound is attained, that is, there exists a suitable choice of $P$ such that for every realisation $y$ we have $\mu_{P,\textup{pos}}(y)=\hat{\mu}_{\textup{pos}}(y)$. The proof is given in Section B.3.
@@
In the finite-dimensional setting, it is shown in [39, Corollary 3.2] that the posterior covariance corresponding to the model (29) agrees with the solution of Problem 4.1. Proposition 7.1 generalises this to infinite dimensions and adds an analogous statement for the posterior mean of model (29): the exact posterior mean of the projected problem (29) with $P$ as in Proposition 7.1 is equal to the optimal low-rank structure-ignoring posterior mean approximation given by Theorem 5.10.
@@ 8 Examples
In this section we consider two typical ill-posed inverse problems to illustrate the proposed framework. We identify the prior-preconditioned Hessian (9a) and its non-self adjoint square root (20) in terms of the functions occurring in the forward problem and the prior. After discretising these expressions, matrix-free methods such as Krylov or Lanczos algorithms and randomized parallel schemes can be used to efficiently approximate the corresponding truncated rank-r SVD; see e.g. [20, 6, 34]. With the $r$ leading eigendirections, the optimal projector $P_{r}$ in Proposition 7.1 can then be constructed, yielding the projected Bayesian inverse problem (29) which contains the essential posterior information. Further details and explanations are provided in Appendix C.
@@
$\partial_{t}u-\partial_{xx}u$ $=0,$ $\text{in }(0,1)\times(0,T),$
$u(\cdot,0)$ $=x^{\dagger},\quad$ $\text{on }(0,1),$
$u(0,\cdot)=u(1,\cdot)$ $=0,$ $\text{on }(0,T].$
The true initial state $x^{\dagger}$ is unknown and needs to be estimated from noisy observations of $u$ at $(x_{i},t_{i})_{i=1}^{n}\subset(0,1)\times(0,T]$. We assume i.i.d. standard Gaussian noise. This problem is similar to [40, Example 3.5] and [20, Section 4.2]. However, in this example we do not observe the entire spatial temperature profile, but observe at finitely many fixed spatial locations, and we consider periodic instead of Dirichlet boundary conditions.
@@ 9 Conclusion
In this work, we have provided the formulation and solution of the low-rank posterior mean approximation problem directly on infinite-dimensional separable Hilbert spaces. We considered approximations that ignore and preserve the structure of the prior-to-posterior mean update in Theorem 5.10 and Theorem 5.11 respectively. To quantify the posterior mean approximation quality, we have considered various loss classes. These loss classes consist of divergences between the exact Gaussian posterior and the approximate Gaussian posterior given by an approximate posterior mean and the exact posterior covariance, after averaging over the data distribution. The chosen divergences are the Hellinger distance and the Renyi, Amari, and forward and reverse KL divergences. These loss classes form a natural extension of the Bayes risk used in finite dimensions in [39], and were used to assess optimality for low-rank approximations to the posterior covariance in [10].
@@ A.1 Hilbert spaces and bounded operators
Suppose that $(h_{n})_{n}\subset\mathcal{H}$ is weakly convergent with limit $h\in\mathcal{H}$, i.e. $\langle h_{n},k\rangle\rightarrow\langle h,k\rangle$ for all $k\in\mathcal{H}$ as $n\rightarrow\infty$. In particular, $\langle Ah,h-h_{n}\rangle\rightarrow 0$. Since the sequence $(\langle h_{n},k\rangle)_{n}$ is bounded for each $k\in\mathcal{H}$, the principle of uniform boundedness, c.f. [11, Theorem III.14.3], implies that $(h_{n})_{n}$ is a bounded sequence. By [33, Theorem VI.11], $(Ah_{n})_{n}$ converges in norm to $Ah$ since $A$ is compact. Thus, $\lvert\langle A(h-h_{n}),h_{n}\rangle\rvert\leq\lVert A(h-h_{n})\rVert\sup_{n}\lVert h_{n}\rVert\rightarrow 0$. We conclude that $\lvert\langle Ah,h\rangle-\langle Ah_{n},h_{n}\rangle\rvert\leq\lvert\langle A(h-h_{n}),h_{n}\rangle\rvert+\lvert\langle Ah,h-h_{n}\rangle\rvert\rightarrow 0$. ∎
@@ A.2 Unbounded operators
[…]
@@ B.1 Proofs of Section 3
By Lemma A.11(i), $(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\in\mathcal{B}(\mathcal{H})$ satisfies $(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}w_{i}=\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$. The above display thus shows $I-\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}=\sum_{i}-\lambda_{i}w_{i}\otimes w_{i}$. This is a nonnegative and compact operator, since $(-\lambda_{i})_{i}\in\ell^{2}([0,1))$. Applying [23, eq. (4.13)] to this operator, we get for any subspace $V_{r}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ of dimension $r$,
@@
where in the final step we use that the ratio $\frac{\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,z\rangle)}{\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,z\rangle)}$ is invariant under scaling of $X$. It remains to show that the final infimum above is attained. Since $\{z\in\mathcal{H}:\ \lVert z\rVert\leq 1\}$ is weakly compact by [11, Theorem V.4.2], the closed subspace $(\mathcal{C}_{\textup{pr}}^{-1/2}V_{r})^{\perp}\cap\{z\in\mathcal{H}:\ \lVert z\rVert=1\}$ of $\{z\in\mathcal{H}:\ \lVert z\rVert\leq 1\}$ is also weakly compact. Furthermore, $\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,z\rangle)=\langle\mathcal{C}_{\textup{pos}}z,z\rangle$ by Lemma A.15, which is weakly continuous in $z$ by Lemma A.9. Similarly, $\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,z\rangle)$ is weakly continuous. Thus the ratio $\frac{\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,z\rangle)}{\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,z\rangle)}$ is weakly continuous on the weakly compact set $(\mathcal{C}_{\textup{pr}}^{-1/2}V_{r})^{\top}\cap\{z\in\mathcal{H}:\ \lVert z\rVert=1\}$. It follows that the infima above are attained, proving (11). ∎
@@ B.2 Proofs of Section 5
This shows that $\mathcal{C}_{\textup{pos}}=S_{\textup{pos}}S_{\textup{pos}}^{*}$, which proves item (i). Item (ii) now immediately follows from [17, Corollary B.3] and the equality $\operatorname{ran}{S_{\textup{pos}}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}.$ For item (iii), we note that by (22a) we have for $h\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, $\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}h=\sum_{i}(1+\lambda_{i})^{1/2}\langle h,w_{i}\rangle w_{i}=h-\sum_{i=1}^{n}\langle h,w_{i}\rangle w_{i}+\sum_{i=1}^{n}(1+\lambda_{i})^{1/2}\langle h,w_{i}\rangle w_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ as a sum of elements of $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, because $(w_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ by Proposition 3.4. Furthermore, if $k\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, then $h\coloneqq\sum_{i}(1+\lambda_{i})^{-1/2}\langle k,w_{i}\rangle w_{i}$ satisfies $h=k-\sum_{i=1}^{n}\langle k,w_{i}\rangle w_{i}+\sum_{i=1}^{n}(1+\lambda_{i})^{-1/2}\langle k,w_{i}\rangle w_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. By (22a), we have $\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}h=\sum_{i}(1+\lambda_{i})^{1/2}\langle h,w_{i}\rangle w_{i}=\sum_{i}\langle k,w_{i}\rangle w_{i}=k$. We conclude that $\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}$ maps $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ onto $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, so that